OpenStudy (timaashorty):

Help please ! In triangle ABC, C is a right angle. If B= 31 degrees and C =14, what are the measures of the remaining sides and angles? round answers to the nearest hundredth.

OpenStudy (anonymous):

i think we need a picture are B and C angles? that is not possible maybe one is an angle and one is a side? or are they both sides? need some more info here

OpenStudy (timaashorty):

C= 14 is a side. C is a right angle.

OpenStudy (anonymous):



OpenStudy (anonymous):

like that?

OpenStudy (timaashorty):

I believe so I was just trying to draw it [:

OpenStudy (anonymous):

ok let me label a couple things

OpenStudy (anonymous):



OpenStudy (anonymous):

then \[\sin(31)=\frac{b}{14}\] so \[b=14\sin(31)\]you will need a calculator to compute that

OpenStudy (anonymous):

also \[\cos(31)=\frac{a}{14}\]so \[a=14\cos(31)\] again a calculator

OpenStudy (timaashorty):

So both A and B will equall = 12?

OpenStudy (timaashorty):

oops is see my mistake lol

OpenStudy (anonymous):

not likely no

OpenStudy (timaashorty):

b= 7.2 a= 12

OpenStudy (anonymous):

\[14\sin(31)=7.2\] about http://www.wolframalpha.com/input/?i=14sin%2831%29

OpenStudy (anonymous):

yeah that looks good \[b=7.2,a=12\]

OpenStudy (timaashorty):

So those are the sides? how would I do the angles?

OpenStudy (timaashorty):

There's still one more step.. How do I find Angle A @sattellite73?

OpenStudy (anonymous):

the angle of any right triangle have to add up to 180

OpenStudy (anonymous):

i meant ANY TRIANGLE (doesn’t have to be a right triangle)

OpenStudy (timaashorty):

180- 90 - 31 ?

OpenStudy (anonymous):

to find your missing angle, compute \[180-90-31\]

OpenStudy (anonymous):

yup

OpenStudy (timaashorty):

59 degrees. Thank you [:!

OpenStudy (anonymous):

yw